# Near-Extremal Black Holes in Modified Gravity via Spectral Methods

### v1

@@ Abstract
Thanks to the LIGO-Virgo-KAGRA collaboration and the Event Horizon Telescope, we can now probe the extreme-gravity dynamics of compact objects through observations of gravitational wave and electromagnetic events. Gravitational waves emitted by binary black hole mergers are generated in the vicinity of their event horizons, and electromagnetic emissions from accretion disks probe regions close to the innermost stable circular orbit of supermassive black holes Abramowicz and Fragile (2013). These signals grant us a magnifying glass with which to test general relativity in the high-curvature regime of astrophysical black holes and to search for deviations predicted by modified gravity at those scales. However, computing theory-specific observables remains a major challenge. Such theories typically involve higher-derivative corrections to the Einstein equations Bueno and Cano (2016); Burger et al. (2020); Dehghani et al. (2012) and may include additional dynamical scalar or vector fields Cano and Ruipérez (2019); Kanti and Tamvakis (1995); Jackiw and Pi (2003a), leading to a system of coupled, nonlinear, elliptic partial differential equations. The absence of exact black hole solutions in these theories significantly hinders our ability to confront them with observational data.
@@
Significant efforts have been put forth to construct rotating black holes in modified gravity theories, inspired by effective-field theory. For example, numerical solutions were computed in scalar Gauss-Bonnet gravity Delgado et al. (2020); Kleihaus et al. (2011, 2016), dynamical Chern-Simons gravity Sullivan et al. (2020, 2021), cubic gravity Burger et al. (2020), and quartic gravity Dehghani et al. (2012) using finite difference methods. Due to their nature, however, these numerically-approximate solutions are not expressed in terms of analytic, closed-form functions, requiring a large number of spatial grid points to achieve sufficient precision for smooth interpolation (especially when in need for their derivatives). Analytic, closed-form, approximate solutions have only been found using the small-coupling and slow-rotation expansion, which enable a perturbative treatment in both the coupling constant and spin parameter $a$ Yunes and Pretorius (2009); Yagi et al. (2012); Maselli et al. (2015); Cano and Ruipérez (2019); Cardoso et al. (2018); Cano et al. (2022). This scheme greatly reduces the complexity of the modified field equations, thus allowing for closed-form solutions, but many astrophysical black holes are observed to have moderate or large spins, rendering “series-in-$a$ solutions” insufficient Abbott et al. (2023, 2025); Reynolds (2013). While, in principle, this problem could be eliminated by extending the spin expansion to a higher order (assuming the series-in-$a$ expansion is absolutely convergent), the computational cost grows rapidly, making such an approach impractical.
@@
To overcome all of these limitations, we here develop a new framework, based on spectral and pseudospectral collocation methods, to construct analytic and closed-form representations of (highly-accurate, approximate) black hole solutions with any spin in modified gravity Fernandes and Mulryne (2022); Liu et al. (2025). The framework consists of first representing the spacetime as a linear, beyond-Einstein deformation of a Kerr black hole, and then peeling off certain analytic behavior from the deformations to ensure certain boundary conditions are satisfied. The metric deformations are then represented as a spectral expansion, which allows the field equations to be decomposed into a set of algebraic equations. For a given spin value, these algebraic equations can be solved by standard matrix techniques to obtain the spectral expansion coefficients. Repeating this approach for many spin values, one obtains the spectral expansion coefficients as a discrete function of spin, which can then be represented as a polynomial in the logarithm of the extremality parameter $\delta\equiv 1-a^{2}$. Such a framework never requires an expansion in small spin or the numerical evaluation of spectral coefficients at every spacetime grid point, instead providing an analytic, closed-form expression (in a given Boyer-Lindquist-like coordinate system) for every component of the metric that is accurate for any spin value Lam et al. (2025).
@@
In this letter, we present this general framework and exemplify it by constructing, for the first time, rotating black hole solutions of dimensionless spin up to 0.999 in scalar Gauss-Bonnet gravity, dynamical Chern-Simons gravity, and axi-dilaton gravity, all of which emerge as low-energy limits of string theory and quantum gravity candidates Maeda et al. (2009); Moura and Schiappa (2007); Cano and Ruipérez (2019, 2022); Alexander and Yunes (2009); Taveras and Yunes (2008). These new “spectral black hole solutions” represent a tremendous improvement in accuracy over previous perturbative solution, especially when considering rapidly-spinning black holes, thus allowing us to investigate various astrophysical observables.
@@
where $R$ is the Ricci scalar associated with the metric $g_{\mu\nu}$ with signature (,+,+,+)(-,+,+,+), $\varphi$ is a dynamical scalar or pseudoscalar field, $\mathscr{L}_{\rm MG}$ is the Lagrangian density modification to the Einstein-Hilbert action [^1] In principle, dynamical vector fields can be included in $\mathscr{L}_{\rm MG}$, but here we focus on scalar-tensor modifications of gravity., and we henceforth employ units in which $G=1=c$. A general feature of modified gravity theories that are motivated by effective-field theory considerations is that $\mathscr{L}_{\rm MG}$ can be represented as a series in curvature invariants that may or may not be non-minimally coupled to dynamical scalar fields Maeda et al. (2009); Moura and Schiappa (2007); Cano and Ruipérez (2019, 2022); Alexander and Yunes (2009); Taveras and Yunes (2008). One can then write $\mathscr{L}_{\rm MG}=\mathscr{L}_{\rm quad}+\mathscr{L}_{\rm cubic}+\ldots$, to represent quadratic and cubic curvature terms, such as Yunes and Stein (2011); Cano and Ruipérez (2019); Yagi et al. (2016); Jackiw and Pi (2003b); Bueno and Cano (2016)
@@
Let us now represent the metric deformation through spectral expansions. We begin by projecting $h_{\alpha\beta}(r,\chi)$ onto an orthogonal spectral basis $u_{n\ell}(r,\chi)$,
@@
Three examples.— To exemplify our spectral framework and demonstrate its validity, we apply it to three specific cases of quadratic gravity: scalar Gauss-Bonnet gravity, dynamical Chern-Simons gravity, and axi-dilaton gravity. The Lagrangian density is that of Eq. (2), with ${\cal{L}}_{\rm MG}={\cal{L}}_{\rm quad}$, and with $V=0$ (massless scalars). In scalar Gauss-Bonnet and in dynamical Chern-Simons gravity, $(\nabla\varphi)^{2}=(\nabla\varphi_{\rm sGB,dCS})^{2}$, $f(\varphi)=\varphi_{\rm sGB,dCS}$ (shift-symmetry), and $\mathscr{Q}=\mathscr{Q}_{\rm sGB,dCS}$, respectively, where $\mathscr{Q}_{\rm sGB}=R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}-4R_{\mu\nu}R^{\mu\nu}+R^{2}$ is the Gauss-Bonnet invariant Yagi et al. (2016); Nair et al. (2019) and $\mathscr{Q}_{\rm dCS}=\tilde{R}^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$ is the Pontryagin invariant Alexander and Yunes (2009); Jackiw and Pi (2003b). In axi-dilaton gravity, $(\nabla\varphi)^{2}=(\nabla\varphi_{\rm sGB})^{2}+(\nabla\varphi_{\rm dCS})^{2}$ and $f(\varphi)\mathscr{Q}=\varphi_{\rm GB}\mathscr{Q}_{\rm sGB}+\varphi_{\rm dCS}\mathscr{Q}_{\rm dCS}$. These coupling functions $f(\varphi)$ can be thought of as a leading-order, small-field approximation of a more sophisticated $f(\varphi)$, when $\mathscr{Q}$ is a topological invariant Chung and Yunes (2024a). Given these Lagrangian densities, one can easily derive the field equations, and the source tensor $\mathscr{S}_{\mu\nu}$ can be computed explicitly (see e.g. Cano and Ruipérez (2019); Yunes and Pretorius (2009)).
@@
Future directions.— The specific black hole solutions obtained here in scalar Gauss-Bonnet, dynamical Chern-Simons and axi-dilaton gravity can now be used to compute astrophysical observables associated with rapidly-rotating black holes, without incurring systematic uncertainties due to finite-difference errors or the slow-rotation expansion. Two key examples are the modeling of quasi-normal modes and extreme mass-ratio inspirals. In both cases, one can use the analytic solutions obtained here as a background on which to study linear vacuum perturbations (to obtain the quasi-normal spectrum, see e.g. Chung and Yunes (2024b, a); Chung et al. (2025)) or matter perturbations (to obtain extreme mass-ratio inspiral waveforms). The former are particularly important to carry out ringdown tests of general relativity with gravitational waves detected by advanced LIGO and its partners, in a way that is robust to systematics, as described e.g. in Chung and Yunes (2025). The latter are similarly important to carry out tests of general relativity with extreme-mass-ratio inspirals using LISA observations Gair et al. (2017); Chamberlain and Yunes (2017); Babak et al. (2017); Maselli et al. (2020); Barsanti et al. (2023).
@@
Acknowledgments.— The authors acknowledge the support from the Simons Foundation through Award No. 896696, the NSF through Grant No. PHY-2207650 and NASA through Grant No. 80NSSC22K0806. The calculations and results reported in this Letter were produced using the computational resources of the Illinois Campus Cluster, a computing resource that is operated by the Illinois Campus Cluster Program (ICCP) in conjunction with National Center for Supercomputing Applications (NCSA), and is supported by funds from the University of Illinois at Urbana-Champaign. The author would like to specially thank the investors of the IlliniComputes initiatives and GravityTheory computational nodes for permitting the authors to execute runs related to this work using the relevant computational resources.

### v2

# Spinning Black Holes in Modified Gravity via Spectral Methods

@@ Abstract
Thanks to the LIGO-Virgo-KAGRA collaboration and the Event Horizon Telescope, we can now probe the extreme-gravity dynamics of compact objects through observations of gravitational wave and electromagnetic events. Gravitational waves emitted by binary black hole mergers are generated in the vicinity of their event horizons, and electromagnetic emissions from accretion disks probe regions close to the innermost stable circular orbit of supermassive black holes Abramowicz and Fragile (2013). These signals grant us a magnifying glass with which to test general relativity in the high-curvature regime of astrophysical black holes and to search for deviations predicted by modified gravity at those scales. However, computing theory-specific observables remains a major challenge. Such theories typically involve higher-derivative corrections to the Einstein equations Bueno and Cano (2016); Burger et al. (2020); Dehghani et al. (2012) and may include additional dynamical scalar or vector fields Cano and Ruipérez (2019); Kanti and Tamvakis (1995); Jackiw and Pi (2003b), leading to a system of coupled, nonlinear, elliptic partial differential equations. The absence of exact black hole solutions in these theories significantly hinders our ability to confront them with observational data.
@@
Significant efforts have been put forth to construct rotating black holes in modified gravity theories, inspired by effective-field theory. For example, numerical solutions were computed in scalar Gauss-Bonnet gravity Delgado et al. (2020); Kleihaus et al. (2011, 2016), dynamical Chern-Simons gravity Sullivan et al. (2020, 2021), cubic gravity Burger et al. (2020), and quartic gravity Dehghani et al. (2012) using finite difference methods. Due to their nature, however, these numerically-approximate solutions are not expressed in terms of analytic, closed-form functions, requiring a large number of spatial grid points to achieve sufficient precision for smooth interpolation (especially when in need for their derivatives). Analytic, closed-form, approximate solutions have only been found using the small-coupling and slow-rotation expansion, which enable a perturbative treatment in both the coupling constant and spin parameter $a$ Yunes and Pretorius (2009); Yagi et al. (2012); Maselli et al. (2015); Cano and Ruipérez (2019); Cardoso et al. (2018); Cano et al. (2022). This scheme greatly reduces the complexity of the modified field equations, thus allowing for closed-form solutions, but many astrophysical black holes are observed to have moderate or large spins, rendering “series-in-$a$ solutions” insufficient Abbott and others (2023); Abac and others (2025); Reynolds (2013). While, in principle, this problem could be eliminated by extending the spin expansion to a higher order (assuming the series-in-$a$ expansion is absolutely convergent), the computational cost grows rapidly, making such an approach impractical.
@@
To overcome all of these limitations, we here develop a new framework, based on spectral and pseudospectral (collocation) methods, to construct analytic and closed-form representations of highly-accurate, approximate black hole solutions with any spin in modified gravity Fernandes and Mulryne (2022); Liu et al. (2025). Our spectral framework consists of first representing the spacetime as a linear, beyond-Einstein deformation of a Kerr black hole [^1] EFTs require solutions to be a smooth deformation of GR solutions. This implies that nonperturbative solutions, such as scalarized black-holes, could not be constructed with our methodology., and then peeling off certain analytic behavior from the deformations to ensure certain boundary conditions are satisfied. The metric deformations are then represented as spectral expansions, which allow the field equations to be decomposed into a set of algebraic equations. For a given spin value, these algebraic equations can be solved by standard matrix techniques to obtain the spectral expansion coefficients. Repeating this approach for many spin values, we obtain the spectral expansion coefficients as a discrete function of spin, which can then be represented as a polynomial in the logarithm of the extremality parameter $\delta\equiv 1-a^{2}$. Such a framework never requires an expansion in small spin or the numerical evaluation of spectral coefficients at every spacetime grid point, instead providing an analytic, closed-form expression (in a given Boyer-Lindquist-like coordinate system) for every component of the metric that is accurate for any spin value Lam et al. (2026).
@@
In this letter, we present this general framework and exemplify it by constructing, for the first time, rotating black hole solutions of dimensionless spin up to 0.999 in scalar Gauss-Bonnet gravity, dynamical Chern-Simons gravity, and axi-dilaton gravity, all of which emerge as low-energy limits of string theory and quantum gravity candidates Maeda et al. (2009); Moura and Schiappa (2007); Cano and Ruipérez (2019, 2022); Alexander and Yunes (2009); Taveras and Yunes (2008). These new “spectral black hole solutions” show orders-of-magnitude improvement in spin in comparison with the previous series-in-$a$ solution, especially when considering rapidly-spinning black holes (see Fig. 2 and Lam et al. (2026)). Moreover, the spectral solutions are analytic in nature, thus allowing us to investigate various astrophysical observables, e.g. ringdown spectra, in modified gravity theories without being subjected to interpolation and other numerical error.
@@
where $R$ is the Ricci scalar associated with the metric $g_{\mu\nu}$ with signature (,+,+,+)(-,+,+,+), $\varphi$ is a dynamical scalar or pseudoscalar field, $\mathscr{L}_{\rm MG}$ is the Lagrangian density modification to the Einstein-Hilbert action [^2] In principle, dynamical vector fields can be included in $\mathscr{L}_{\rm MG}$, but here we focus on scalar-tensor modifications of gravity., and we henceforth employ units in which $G=1=c$. A general feature of modified gravity theories that are motivated by effective-field theory considerations is that $\mathscr{L}_{\rm MG}$ can be represented as a series in curvature invariants that may or may not be non-minimally coupled to dynamical scalar fields Maeda et al. (2009); Moura and Schiappa (2007); Cano and Ruipérez (2019, 2022); Alexander and Yunes (2009); Taveras and Yunes (2008). One can then write $\mathscr{L}_{\rm MG}=\mathscr{L}_{\rm quad}+\mathscr{L}_{\rm cubic}+\ldots$, to represent quadratic and cubic curvature terms, such as Yunes and Stein (2011); Cano and Ruipérez (2019); Yagi et al. (2016); Jackiw and Pi (2003a); Bueno and Cano (2016)
@@
Let us now represent the metric deformation through spectral expansions. Assuming $h_{\alpha\beta}$ and its derivatives are finite and smooth outside of the event horizon, we begin by projecting $h_{\alpha\beta}(r,\chi)$ onto an orthogonal spectral basis $u_{n\ell}(r,\chi)$,
@@
Three examples.— To exemplify our spectral framework and demonstrate its validity, we apply it to three specific cases of quadratic gravity: scalar Gauss-Bonnet gravity, dynamical Chern-Simons gravity, and axi-dilaton gravity. The Lagrangian density of these theories are that of Eq. (2), with ${\cal{L}}_{\rm MG}={\cal{L}}_{\rm quad}$, and with $V=0$ (massless scalars). The coupling constants of these theories in ${\cal L}_{\rm MG}$ are observed to be small Chung and Yunes (2025); Lyu et al. (2022); Silva et al. (2023, 2020), and thus, it is valid to apply this perturbative-in-$\zeta$ framework. In scalar Gauss-Bonnet and in dynamical Chern-Simons gravity, $(\nabla\varphi)^{2}=(\nabla\varphi_{\rm sGB,dCS})^{2}$, $f(\varphi)=\varphi_{\rm sGB,dCS}$ (shift-symmetry), and $\mathscr{Q}=\mathscr{Q}_{\rm sGB,dCS}$, respectively, where $\mathscr{Q}_{\rm sGB}=R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}-4R_{\mu\nu}R^{\mu\nu}+R^{2}$ is the Gauss-Bonnet invariant Yagi et al. (2016); Nair et al. (2019) and $\mathscr{Q}_{\rm dCS}=\tilde{R}^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$ is the Pontryagin invariant Alexander and Yunes (2009); Jackiw and Pi (2003a). In axi-dilaton gravity, $(\nabla\varphi)^{2}=(\nabla\varphi_{\rm sGB})^{2}+(\nabla\varphi_{\rm dCS})^{2}$ and $f(\varphi)\mathscr{Q}=\varphi_{\rm GB}\mathscr{Q}_{\rm sGB}+\varphi_{\rm dCS}\mathscr{Q}_{\rm dCS}$. These coupling functions $f(\varphi)$ can be thought of as a leading-order, small-field approximation of a more sophisticated $f(\varphi)$, when $\mathscr{Q}$ is a topological invariant Chung and Yunes (2024a). Given these Lagrangian densities, one can easily derive the field equations, and the source tensor $\mathscr{S}_{\mu\nu}$ can be computed explicitly (see e.g. Cano and Ruipérez (2019); Yunes and Pretorius (2009)).
@@
Future directions.— The specific black hole solutions obtained here in scalar Gauss-Bonnet, dynamical Chern-Simons and axi-dilaton gravity can now be used to compute astrophysical observables associated with rapidly-rotating black holes, without incurring uncertainties due to finite-difference errors (especially from derivatives of the metric) or the slow-rotation expansion. Two key examples are the modeling of quasi-normal modes and extreme mass-ratio inspirals. In both cases, one can use the analytic solutions obtained here as a background on which to study linear vacuum perturbations (to obtain the quasi-normal spectrum, see e.g. Chung and Yunes (2024b, a); Chung et al. (2025)) or matter perturbations (to obtain extreme mass-ratio inspiral waveforms). The former are particularly important to carry out ringdown tests of general relativity with gravitational waves detected by advanced LIGO and its partners, in a way that is robust to systematics, as described e.g. in Chung and Yunes (2025). The latter are similarly important to carry out tests of general relativity with extreme-mass-ratio inspirals using LISA observations Gair et al. (2017); Chamberlain and Yunes (2017); Babak et al. (2017); Maselli et al. (2020); Barsanti et al. (2023).
@@
Acknowledgments.— The authors acknowledge the support from the Simons Foundation through Award No. 896696, the Simons Foundation International under grant SFI-MPS-BH-00012593-01, the NSF through Grant No. PHY-2207650 and PHY 25-12423, and NASA through Grant No. 80NSSC22K0806. A. K. W. C also acknowledge the Herchel Smith Fellowship at the University of Cambridge for partial support of this work. The calculations and results reported in this Letter were produced using the computational resources of the Illinois Campus Cluster, a computing resource that is operated by the Illinois Campus Cluster Program (ICCP) in conjunction with National Center for Supercomputing Applications (NCSA), and is supported by funds from the University of Illinois at Urbana-Champaign. The author would like to specially thank the investors of the IlliniComputes initiatives and GravityTheory computational nodes for permitting the authors to execute runs related to this work using the relevant computational resources.